# Efficient Visual State Space Model for Image Deblurring

### v1

@@ 4.1 Datasets and implementation
Following existing state-of-the-art methods, we evaluate our approach on the commonly used GoPro dataset by Nah et al. [31], HIDE dataset by Shen et al. [36], and RealBlur dataset by Rim et al. [35]. The GoPro dataset contains $2103$ images for training and $1111$ images for testing. The HIDE dataset includes $2025$ images mainly about humans for testing. The RealBlur dataset [35] contains RealBlur-J and RealBlur-R subsets generated by different post-processing strategies. It uses $182$ scenes for training and $50$ scenes for testing. For fair comparisons, we follow the protocols of these datasets to evaluate our method.
@@
The shallow feature $F_{s}$ has a channel number of 48 and the numbers of EVSS modules in the encoder/decoder from level-$1$ to level-$3$ are [6, 6, 12]. We use the ADAM optimizer [20] with default parameters in the training process. We use the data augmentation method with the flipping and rotation operations to generate training data. We apply the progressive training similar to but simpler than [54]: the training starts with the patch size of $128\times 128$ pixels and batch size $64$ for $300,000$ iterations, where the learning rate gradually reduces from $1\times 10^{-3}$ to $1\times 10^{-7}$. Then the patch size is enlarged to $256\times 256$ pixels with $16$ batches for $300,000$ iterations where the learning rate is initialized as $5\times 10^{-4}$ and decreases to $1\times 10^{-7}$. The learning rate is updated based on the Cosine Annealing scheme. Unless otherwise specified, all experiments are conducted with the PyTorch framework on NVIDIA RTX 4090 GPUs. The training code and test model are available at https://github.com/kkkls/EVSSM.
Table 1: Quantitative evaluations of the proposed method against state-of-the-art ones on the GoPro dataset [31] in terms of PSNR and SSIM.
Method SRN [40] DMPHN [55] MIMO-UNet+ [5] MPRNet [53] MAXIM [43] Uformer [48] Restormer [54]
PSNR (dB) 30.26 31.20 32.45 32.66 32.86 33.06 32.92
SSIM 0.9342 0.9453 0.9567 0.9589 0.9616 0.9670 0.9611
Method Stripformer [42] Restormer-local [6] NAFNet [3] GRL [25] FFTformer [21] CU-mamba [7] EVSSM (ours)
PSNR (dB) 33.08 33.57 33.71 33.93 34.21 33.53 34.50
SSIM 0.9624 0.9656 0.9668 0.9680 0.9692 0.9650 0.9712
Table 2: Quantitative evaluations of the proposed method against state-of-the-art ones on the real-world dataset [35] in terms of PSNR and SSIM.
Dataset Method DeblurGAN-v2 [23] SRN [40] MIMO-Unet+ [5] BANet [41] DeepRFT+ [30] Stripformer [42] FFTformer [21] EVSSM (ours)
RealBlur-J PSNR (dB) 29.69 31.38 31.92 32.00 32.19 32.48 32.62 34.15
SSIM 0.8703 0.9091 0.9190 0.9230 0.9305 0.9290 0.9326 0.9450
RealBlur-R PSNR (dB) 36.44 38.65 - 39.55 39.84 39.94 40.11 41.04
SSIM 0.9347 0.9652 - 0.9710 0.9721 0.9739 0.9732 0.9770
Table 3: Quantitative evaluations of the proposed method against state-of-the-art ones on the HIDE dataset [36] in terms of PSNR and SSIM.
Method SRN [40] DMPHN [55] SAPHN [39] MIMO-UNet+ [5] MPRNet [53] Uformer [48] MPRNet-local [6]
PSNR (dB) 28.36 29.09 29.98 29.99 30.96 30.90 31.19
SSIM 0.9040 0.9240 0.9300 0.9304 0.9397 0.9530 0.9418
Method Restormer [54] Stripformer [42] Restormer-local [6] NAFNet [3] GRL [25] FFTformer [21] EVSSM (ours)
PSNR (dB) 31.22 31.03 31.49 31.31 31.65 31.62 31.97
SSIM 0.9423 0.9395 0.9447 0.9427 0.947 0.9455 0.9501
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
(e) Restormer-local [6] (f) NAFNet [3] (g) FFTformer [21] (h) EVSSM (ours)
@@ 4.2 Comparisons with the state of the art
We first evaluate the performance of the proposed approach on the GoPro dataset [31]. We compare our method against state-of-the-art approaches, including CNN-based ones (SRN [40], DMPHN [55], MIMO-UNet+ [5], MPRNet [53], NAFNet [3]), Transformer-based ones (Uformer [48], Restormer [54], Stripformer [42], Restormer-local [6], GRL [25], and FFTformer [21]), MLP-based ones (MAXIM [43]), and SSM-based ones (CU-mamba [7]). We retrain or fine-tune the deep learning-based methods for fair comparisons if they are not trained on the benchmarks. As CU-mamba [7] does not provide training and test codes, for fair comparison, we compare our result with the reported result of [7]. We use the PSNR and SSIM as the evaluation metrics to measure the quality of each restored image. Table 1 shows the quantitative evaluation results, where our approach outperforms other methods with the highest PSNR and SSIM values.
Figure 2 shows evaluation results on the GoPro dataset [31]. Since CNN-based methods cannot effectively utilize global information, the images restored by [5, 3] still contain severe blur residuals as shown in Figure 2(c) and (f). Although Transformer-based methods can model the global context, the methods [42, 6, 21] have adopted various approximations to reduce the computational cost. This affects their ability to model the global information, and as a result, some main structures, e.g., tire, and wheel, are not restored well (see Figure 2(d), (e) and (g)). In contrast to existing Transformer-based methods, we propose a simple yet effective visual state space model for image deblurring, which is effective at exploring non-local information with low computational costs. As shown in Figure 2(h), our approach generates better results, where the text and tire are much clearer.
@@
Using the same protocols, we evaluate our approach on the real-world blurry dataset by Rim et al. [35]. Table 2 shows that the proposed method outperforms previous work significantly, improving the PSNR by at least $1.53$dB and $0.93$dB on the datasets of RealBlur-J and RealBlur-R, respectively. Figure 3 shows the visual comparisons on the RealBlur dataset, where our method generates the result with clearer characters and finer structural details.
Refer to caption Refer to caption Refer to caption
(a) Blurred input (b) GT (c) SRN [40] (d) DeblurGAN-v2 [23]
Refer to caption Refer to caption Refer to caption
(e) MIMO-UNet+ [5] (f) Stripformer [42] (g) FFTformer [21] (h) EVSSM (ours)
@@
Method DSC [29] MSPFN [19] RCDNet [45] DualGCN [10] SPDNet [52] Restormer [54] IDT [50] DRSformer [4] EVSSM (ours)
PSNR (dB) 34.95 43.43 43.36 44.18 43.20 47.98 47.35 48.54 48.78
SSIM 0.9416 0.9843 0.9831 0.9902 0.9871 0.9921 0.9930 0.9924 0.9951
Method DCP [18] MSCNN [33] GFN [34] MSBDN [8] PFDN [9] FFA-Net [32] AECRNet [49] DehazeFormer [38] EVSSM (ours)
PSNR (dB) 17.88 20.31 23.52 28.56 28.15 29.96 28.52 31.45 31.66
SSIM 0.816 0.863 0.905 0.966 0.962 0.973 0.964 0.980 0.968
@@
We also evaluate our method on the real-world dataset [46] for image deraining and RESIDE-6K dataset [24] for image dehazing. Tables 4-5 show that the proposed approach performs favorably against state-of-the-art methods on image deraining and dehazing datasets, which demonstrates the generalization ability of our method on other image restoration tasks.
@@ 5 Analysis and Discussion
We have shown that the proposed efficient visual state space model generates favorable results compared to state-of-the-art methods. In this section, we provide more analysis of the proposed method and discuss the effect of the main components. For the ablation studies in this section, we train our approach and all the baselines on the GoPro dataset using a batch size of $64$ and a patch size of $128\times 128$ pixels.
@@
The core of our proposed EVS block is the geometric transformation (4) applied at the beginning of this block. To demonstrate the effect of the EVS block, we first remove the geometric transformation (one-direction for short, which performs scanning in one direction as in [13]) and train this baseline model using the same settings as ours. Table 6 shows that the PSNR of our approach is $0.14$dB higher than this baseline method. The state space model requires flattening the image feature into a one-dimensional sequence, which compromises the spatial structural information of the visual data. Compared to the baseline, our method with the geometric transformation can better explore non-local information. Meanwhile, the number of parameters and the FLOPs of our method are the same as those of the baseline, and their runtimes are also almost identical, which demonstrates the effectiveness of the proposed EVS block in improving the ability of the state space model for handling visual data with a minimal increase in computational costs.
@@
Scanning mode      Parameters (M) FLOPs (G)      Runtime (ms) PSNR/SSIM
one-direction 17.1 126 87.9 33.99/0.9683
two-direction 18.3 135 122.7 34.08/0.9687
four-direction 20.1 148 182.6 34.05/0.9686
EVS (ours) 17.1 126 88.7 34.13/0.9690
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Blurred input (b) one-direction (c) two-direction (d) four-direction (e) EVS (ours)
Figure 4: Effect of the proposed EVS block. Our method with the proposed EVS block is more effective at recovering clearer structures.
In addition, we also compare with two baselines that respectively perform scanning in two directions (two-direction for short) and four directions [27] (four-direction for short). As Table 6 shows, although scanning in multi-directions can alleviate the limitation of the state space model on handling visual data, it leads to an increase in the number of network parameters and the computational complexity, resulting in significantly longer runtimes, cf. $122.7$ms for one-direction vs. $182.6$ms for four-direction vs. $88.7$ms for our approach. Note that the results of scanning in four-directions are slightly lower than those of scanning in two-directions. This is because the methods that scan in multiple directions simply employ the summation followed by normalization to fuse features extracted from different scanning directions. Thus, the multi-direction information is not effectively and fully utilized. In contrast to simultaneously scanning in multiple directions, our approach applies geometric transformations to the input feature at the beginning of each EVS block. This allows each scan to capture the contextual information from different directions and mitigates the increase in computational complexity and runtime. Figure 4 demonstrates that our approach can deblur images better than other methods, where the structures of the windows are restored well (see Figure 4(e)).
@@
Method Scanning in one direction Transpose Flip Parameters (M) FLOPs (G) Runtime (ms) PSNR/SSIM
w/o F&T 17.1 126 87.9 33.99/0.9683
w/o F 17.1 126 88.4 34.05/0.9686
w/o T 17.1 126 88.5 34.03/0.9684
EVSSM (ours) 17.1 126 88.7 34.13/0.9690
@@
Method DMPHN [55] IPT [2] MPRNet-local [6] Restormer-local [6] GRL [25] FFTformer [21] EVSSM (ours)
Parameters (M) 21.7 114 20.1 26.1 20.20 16.6 17.1
FLOPs (G) 234 376 760 155 1289 131 126
Runtime (ms) 143 298 95 286 518 132 89
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Blurred input (b) w/o F&T (c) w/o F (d) w/o T (e) EVSSM (ours)
Figure 5: Effect of the geometric transformations. Compared to the results in (b)-(d), our approach with both the filp and transpose transformations generates a clearer image in (e).
Effectiveness of the geometric transformation.
In the EVS block, we adopt two classical image geometric transformations in the proposed EVS block: flip and transpose. To demonstrate their effectiveness, we individually remove the flip transformation (w/o F for short), the transpose transformation (w/o T for short), and both flip and transpose transformations (w/o F&T for short). The comparison results in Table 7 demonstrate that applying the flip or transpose transformation can achieve better results, improving the PSNR by at least $0.04$dB. Our approach, which uses both flip and transpose transformations, outperforms all these baseline methods without significantly increasing the computational cost or runtime. The visual comparisons in Figure 5 further demonstrate the effectiveness of our approach, where the wheel recovered by our method is much clearer.
@@
We develop an effective and efficient method that explores the the properties of the state space model for high-quality image restoration. However, we have only considered simple transformations such as flip and transpose so far. In future work, we will consider more powerful transformation methods, such as the polar coordinate transformation, to better characterize the spatial information of visual data with SSMs.

### v2

@@ 4.1 Datasets and implementation
Following existing state-of-the-art methods, we evaluate our approach on the commonly used GoPro dataset by Nah et al. [26], HIDE dataset by Shen et al. [31], and RealBlur dataset by Rim et al. [30]. The GoPro dataset contains $2103$ images for training and $1111$ images for testing. The HIDE dataset includes $2025$ images mainly about humans for testing. The RealBlur dataset [30] contains RealBlur-J and RealBlur-R subsets generated by different post-processing strategies. It has $3758$ images for training and $980$ images for testing. For fair comparisons, we follow the protocols of these datasets to evaluate our method.
@@
The shallow feature $F_{s}$ has a channel number of 48 and the numbers of EVSS modules in the encoder/decoder from level-$1$ to level-$3$ are [6, 6, 12]. We use the AdamW optimizer with default parameters in the training process. We use the data augmentation method with the flipping and rotation operations to generate training data. We apply progressive training similar to but simpler than [48]: the training starts with the patch size of $128\times 128$ pixels and the batch size of $64$ for $300,000$ iterations, where the learning rate gradually reduces from $1\times 10^{-3}$ to $1\times 10^{-7}$. Then the patch size is enlarged to $256\times 256$ pixels with $16$ batches for $300,000$ iterations where the learning rate is initialized as $5\times 10^{-4}$ and decreases to $1\times 10^{-7}$. The learning rate is updated based on the Cosine Annealing scheme. Unless otherwise specified, all experiments are conducted with the PyTorch framework on NVIDIA RTX 4090 GPUs. The training code and models are available at EVSSM.
Table 1: Quantitative evaluations on the GoPro dataset [26] in terms of PSNR and SSIM.
GoPro HIDE
Methods PSNR (dB) SSIM PSNR (dB) SSIM
SRN [35] 30.26 0.9342 28.36 0.9040
DMPHN [49] 31.20 0.9453 29.09 0.9240
SAPHN [34] 31.85 0.9480 29.98 0.9300
MIMO-Unet+ [4] 32.45 0.9567 29.99 0.9304
MPRNet [47] 32.66 0.9589 30.96 0.9397
Restormer [48] 32.92 0.9611 31.19 0.9418
Uformer-B [43] 33.06 0.9670 30.90 0.9530
Stripformer [37] 33.08 0.9624 31.03 0.9395
DeepRFT+ [25] 33.23 0.9632 31.42 0.9440
MPRNet-local [5] 33.31 0.9637 31.19 0.9418
Restormer-local [5] 33.57 0.9656 31.49 0.9447
NAFNet [2] 33.71 0.9668 31.31 0.9427
GRL [22] 33.93 0.9680 31.65 0.9470
FFTformer [19] 34.21 0.9692 31.62 0.9455
CU-mamba [6] 33.53 0.9650 31.47 0.9440
MLWNet-B [12] 33.83 0.9676 31.06 0.9320
EVSSM 34.51 0.9713 31.99 0.9503
Table 2: Quantitative evaluations on the RealBlur dataset [30] in terms of PSNR and SSIM.
Realblur-R Realblur-J
Methods PSNR (dB) SSIM PSNR (dB) SSIM
DeblurGAN-v2 [20] 36.44 0.9347 29.69 0.8703
SRN [35] 38.65 0.9652 31.38 0.9091
MIMO-Unet+ [4] - - 31.92 0.9190
BANet [36] 39.55 0.9710 32.00 0.9230
DeepRFT+ [25] 39.84 0.9721 32.19 0.9305
Stripformer [37] 39.84 0.9737 32.48 0.9290
FFTformer [19] 40.11 0.9753 32.62 0.9326
MLWNet [12] 40.69 0.9755 33.84 0.9405
EVSSM 41.27 0.9776 34.34 0.9456
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
(e) Restormer-local [5] (f) NAFNet [2] (g) MLWNet [12] (h) EVSSM (ours)
@@ 4.2 Comparisons with the state of the art
We first evaluate the performance of the proposed approach on the GoPro dataset [26]. We compare our method against state-of-the-art approaches, including CNN-based ones (SRN [35], DMPHN [49], MIMO-UNet+ [4], MPRNet [47], NAFNet [2]), Transformer-based ones (Uformer [43], Restormer [48], Stripformer [37], Restormer-local [5], GRL [22], and FFTformer [19]), MLP-based ones (MAXIM [38]), and SSM-based ones (CU-mamba [6]). We retrain or fine-tune the deep learning-based methods for fair comparisons if they are not trained on the benchmarks. As CU-mamba [6] does not provide training and test codes, for fair comparison, we compare with their reported results [6]. We use the PSNR and SSIM as the evaluation metrics to measure the quality of each restored image. Table 1 shows the quantitative evaluation results, where our approach outperforms other methods with the highest PSNR and SSIM values.
Figure 2 shows the evaluation results on the GoPro dataset [26]. Since CNN-based methods cannot effectively utilize global information, the images restored by [4, 2] still contain severe blur residuals as shown in Figure 2(c) and (f). Although Transformer-based methods can model the global context, the methods [37, 5, 19] have adopted various approximations to reduce the computational cost. This affects their ability to model the long-range information, and as a result, some main structures, e.g., tire, and wheel, are not restored well (see Figure 2(d) and (e)). In contrast to existing Transformer-based methods, we propose a simple yet effective visual state space model for image deblurring, which is effective at exploring non-local information with low computational costs. As shown in Figure 2(h), our method produces clearer results, especially for text and tire.
@@
Using the same protocols, we evaluate our approach on the real-world blurry dataset by Rim et al. [30]. Table 2 shows that the proposed method outperforms previous work significantly, improving the PSNR by at least $0.58$dB and $0.5$dB on the datasets of RealBlur-J and RealBlur-R, respectively. Figure 3 shows the visual comparisons, where our method generates the result with clearer characters and finer structural details.
Method MSPFN [18] RCDNet [40] DualGCN [9] SPDNet [46]
PSNR (dB) 43.43 43.36 44.18 43.20
SSIM 0.9843 0.9831 0.9902 0.9871
Method Restormer [48] IDT [45] DRSformer [3] EVSSM (ours)
PSNR (dB) 47.98 47.35 48.54 49.00
SSIM 0.9921 0.9930 0.9924 0.9954
Method MSCNN [28] GFN [29] MSBDN [7] PFDN [8]
PSNR (dB) 20.31 23.52 28.56 28.15
SSIM 0.863 0.905 0.966 0.962
Method FFA-Net [27] AECRNet [44] DehazeFormer [33] EVSSM (ours)
PSNR (dB) 29.96 28.52 31.45 32.05
SSIM 0.973 0.964 0.980 0.980
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Blurred input (b) GT (c) SRN [35] (d) MIMO-UNet+ [4]
Refer to caption Refer to caption Refer to caption Refer to caption
(e) Stripformer [37] (f) FFTformer [19] (g) MLWNet [12] (h) EVSSM (ours)
@@
We also evaluate our method on the real-world dataset [42] for image deraining and RESIDE-6K dataset [21] for image dehazing. Tables 3-4 show that the proposed approach performs favorably against state-of-the-art methods on image deraining and dehazing datasets, which demonstrates the generalization ability of our method on other image restoration tasks.
@@ 5 Analysis and Discussion
We have shown that the proposed efficient visual state space model generates favorable results compared to state-of-the-art methods. In this section, we provide more analysis of the proposed method and discuss the effect of the main components. For the ablation studies in this section, we train our approach and all the baselines on the GoPro dataset using a batch size of $64$ and a patch size of $128\times 128$ pixels for 300K iterations with the initial learning rate $1\times 10^{-3}$ gradually reduced to $1\times 10^{-6}$ with the cosine annealing schedule.
@@
The core of our proposed EVS block is the geometric transformation (4) applied at the beginning of this block. To demonstrate the effect of the EVS block, we first remove the geometric transformation (one-direction for short, which performs scanning in one direction as in [13]) and train this baseline model using the same settings as ours. Table 5 shows that the PSNR of our approach is $0.26$dB higher than this baseline method. The state space model requires flattening the image feature into a one-dimensional sequence, which compromises the spatial structural information of the visual data. Compared to the baseline, our method with the geometric transformation can better explore non-local information. Meanwhile, the number of parameters and the FLOPs of our method are the same as those of the baseline, and their runtimes are also almost identical, which demonstrates the effectiveness of the proposed EVS block in improving the ability of the state space model for handling visual data with a minimal increase in computational costs.
@@
Scanning mode Params FLOPs Runtime PSNR (dB)/SSIM
one-direction 17.1M 126G 87.9ms 33.89/0.9671
two-direction 18.3M 135G 122.7ms 33.08/ 0.9687
four-direction 20.1M 148G 182.6ms 33.05/ 0.9686
EVS (ours) 17.1M 126G 88.8ms 34.15/0.9690
In addition, we compare with two baselines that respectively perform scanning in two directions (two-direction for short) and four directions [24] (four-direction for short). As Table 5 shows, although scanning in multi-directions can alleviate the limitation of the state space model on handling visual data, it leads to an increase in the number of network parameters and the computational complexity, resulting in significantly longer runtimes, cf. $122.7$ms for one-direction vs. $182.6$ms for four-direction vs. $88.7$ms for our approach. Note that the results of scanning in four directions are slightly lower than those of scanning in two directions. This is because the methods that scan in multiple directions simply employ the summation followed by normalization to fuse features extracted from different scanning directions. Thus, the multi-directional information is not effectively and fully utilized. In contrast to simultaneously scanning in multiple directions, our approach applies geometric transformations to the input feature at the beginning of each EVS block. This allows each scan to capture contextual information from different directions and mitigates the increase in computational complexity and runtime.
@@
Method Params FLOPs Runtime PSNR (dB)/SSIM
w/o F&T 17.1M 126G 87.9ms 33.89/0.9671
w/o F 17.1M 126G 88.4ms 34.05/0.9685
w/o T 17.1M 126G 88.4ms 34.03/0.9684
EVSSM (ours) 17.1M 126G 88.8ms 34.15/0.9690
Effectiveness of the geometric transformation.
In the EVS block, we adopt two classical image geometric transformations: flip and transpose. To demonstrate their effectiveness, we individually remove the flip transformation (w/o F for short), the transpose transformation (w/o T for short), and both flip and transpose transformations (w/o F&T for short). The comparison results in Table 6 show that applying the flip or transpose transformation can achieve better results, improving the PSNR by at least $0.1$dB. Our approach, which employs both flip and transpose transformations, outperforms all these baseline methods, with a notable improvement in PSNR by 0.26 dB, without significantly increasing the computational cost or runtime.
Effect of the efficient discriminative frequency domain-based FFN (EDFFN).
The EDFFN is mainly used to reduce computational complexity while maintaining performance. To demonstrate its effectiveness, we compare two baseline methods that respectively only employ DFFN (w/ only DFFN for short) and EDFFN (w/ only EDFFN for short) in the encoder/decoder. In contrast to DFFN that applies frequency-domain based operations in the middle of the FFN, where the number of channels of the feature is 6×6$\times$6 × that of the input feature, our approach performs frequency-domain screening on the features at the final stage of the FFN network where the number of channels is the same as that of the input feature. As Table 7 shows, the runtime of the baseline with EDFFN is half that of the baseline with DFFN and the number of model parameters is reduced by 1.3M. Concurrently, there is no loss in the performance of the baseline with EDFFN.
Table 7: Effectiveness of the EDFFN, evaluated on the GoPro dataset [26].
Method Params GPU memory Runtime PSRN (dB)/SSIM
w/ only DFFN 11.5M 12.2G 75ms 33.67/0.9664
w/ only EDFFN 10.2M 8.3G 36ms 33.68/0.9661
@@
Method IPT [1] Restormer-local [5] GRL [22] FFTformer [19] EVSSM
Params 114M 26.1M 20.20M 16.6M 17.1M
FLOPs 376G 155G 1289G 131G 126G
Runtime 298ms 286ms 518ms 132ms 89ms
@@
We develop an effective and efficient method that explores the properties of the state space model for high-quality image restoration. However, we have only considered simple transformations such as flip and transpose so far. In future work, we will consider more powerful transformation methods, such as the polar coordinate transformation, to better characterize the spatial information of visual data with SSMs.
@@ 6 Conclusion
Acknowledgments. This work has been supported in part by the National Natural Science Foundation of China (Nos. 62272233, U22B2049, and 62332010). M.-H. Yang is supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korean Government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project).